2. A o 5-kilogram cube slides at 20 m/s on a smoot
Answer #1

Apply conservation of momentum to the sphere and cube

m1u1 + m2 u2 = m1v1 + m2 v2

The final speed of the sphere after collisiion is

v2 = 2 m1/ m1 + m2 ) u1 + (m2- m1/ m1 + m2) u2

since u2 = 0

v2 = 2 m1/ m1 + m2 ) u1

=2 (0.5)/(0.5+4(0.5) * 20

=8 m/s

the final velocity of the cube

v1 = ( m1- m2/ m1 + m2) u1

= (0.5-4(0.5)/0.5+4(0.5) * 20

=-12 m/s

(a)

Impulse = change in momentum of the sphere

=m2( v2- u2)

= 4(0.5) (8 m/s - 0)

= 16 kg m/s

(b)

change in momentum of the sphere is

=m2( v2- u2)

= 4(0.5) (8 m/s - 0)

= 16 kg m/s

(c)

v2 = 2 m1/ m1 + m2 ) u1

=2 (0.5)/(0.5+4(0.5) * 20

=8 m/s

(d)

v1 = ( m1- m2/ m1 + m2) u1

= (0.5-4(0.5)/0.5+4(0.5) * 20

=-12 m/s

magnitude of velocity is 12 m/s

toward left side